OpenStudy (anonymous):

I WILL GIVE A MEDAL!! Part A: Solve F = C + 32 for C. (4 points) Part B: Determine the value of C when F = 10 °F. (2 points) Part C: Solve -np - 70 > 40 for n. Show your work. (4 points)

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

Connections Academy help and answers https://plus.google.com/communities/102495724813457384957

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

but i was hoping to get help here

OpenStudy (anonymous):

can you help at all?!

OpenStudy (anonymous):

I go to CA as well :) Okay, so. A: F - 32 = C B: 10° F - 32 = -22 C: ugh i cant do c rn im sorry DX

OpenStudy (zarkon):

@helpcomonplzhelpme your conversion formula is not correct

OpenStudy (anonymous):

uhg

OpenStudy (anonymous):

no it is

OpenStudy (anonymous):

I don't think it's meant to be accurate, CA does that sometimes and it's kind of stupid and confusing.

OpenStudy (zarkon):

\[F=\frac{9}{5}C+32\]

OpenStudy (zarkon):

ah...that is not good

OpenStudy (anonymous):

thats what they gave me

OpenStudy (anonymous):

Yeah. >.< I hate it so i just take away the stupid degree stuff and just think of it as variables

OpenStudy (zarkon):

that would be better

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i got a

OpenStudy (anonymous):

all i need is b and c

OpenStudy (anonymous):

I told you B but if you still need C I would post it as a new, separate question.

OpenStudy (anonymous):

grrrr

OpenStudy (anonymous):

B is wrong

OpenStudy (anonymous):

Sorry :/ I can't do inequalities and sh** right now. My brain is to ajumble

OpenStudy (anonymous):

The formulaa is 2f/9 -32=C

OpenStudy (anonymous):

A: F - 32 = C B: 10° F - 32 = -22 Ignore that other thing.

OpenStudy (anonymous):

i don't think that is correct

OpenStudy (anonymous):

It is. Trust meee

OpenStudy (anonymous):

i will make a new question

OpenStudy (anonymous):

Up to you~

OpenStudy (anonymous):

well

OpenStudy (anonymous):

I think you have A wrong which messed up B

OpenStudy (anonymous):

F=95C+32

OpenStudy (anonymous):

not that

OpenStudy (anonymous):

They were both correct. I've done the test you're in before.

OpenStudy (anonymous):

I got C wrong then so idk though.

OpenStudy (anonymous):

what test then?

OpenStudy (anonymous):

I don't remember I did it like 2 years ago

OpenStudy (anonymous):

If you don't trust me then whatever i dont care

OpenStudy (anonymous):

sorry its just i need to get this